NASA Space Station On-Orbit Status 9 September 2004

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Thursday, September 9, 2004

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.

  Update on Elektron:  The Russian O2 generator is still off.   [The Elektron, which contains the BZh-7 fluid unit as main component, shut down twice because of increased pressure in the assembly's O2 supply line, not due to air bubbles in the BZh-7.  Moscow identified two possible causes of this, viz. line clogging by crystallization deposits, or hydrogen (H2) having leaked into it.  The former cause is currently the preferred one.  Tomorrow, the crew will perform extensive IFM (in-flight maintenance) on the Elektron, focusing on (a) replacement of the gas analyzer for H2 in the O2 line found faulty (a 2.5 to 3 hrs job), and flushing the (probably) clogged O2 line with water and purging it with nitrogen (N2).]

Service Module Atmosphere Revitalization Subsystem, Book 2, Mission Operations Directorate, 9 October 2000 [Acrobat] According to this document's introduction "This book contains information for the crew about procedures and rules for the atmosphere revitalization subsystem, Elektron, Vozdukh, Micropurification Unit, and Fire Detection and Suppression Subsystem operations, as well as their schematic and operation logic."

Update on cabin atmosphere:  With the Elektron off, cabin pressure and ppO2 (oxygen partial pressure) have been decreasing by roughly 3.5 lb/day (= 2.5 mmHg/day).  O2 margins without taking any actions are good for 10+ days.  But cabin air total pressure will reach the lower limit of 14.2 psi (Flight Rule spec) around the morning of Saturday (9/11).  Having accepted a slight drop below the limit, the ISS Mission Management Team (IMMT) will decide Monday (9/13) on the preferred source of the required repressurization.   [Sources are:  O2 from Progress 15P, N2 from U.S. Airlock HPT (high-pressure tank), and O2 from A/L HPT (least desirable).  There is 62 lbs (28 kg) of O2 in 15P, enough for 16 days.  There are also 84 SFOG (solid-fuel oxygen generator) "candles" on board, which would supply another 42 days' worth of O2.  In addition, the U.S. HPTs contain 150 lb of O2 and 180 lb of N2.  Next O2 supply will arrive with Progress 16 on 11/26.]

Yesterday's exploratory work on the older BZh-5 Fluid Unit for the Elektron was successful, gaining a spare for the current assembly.  [The crew connected a separate UTs circulation unit (providing external micropumps), launched on Soyuz 8S, to the BZh to attempt to remove any bubbles from the unit and recover its functionality.  BZh-5 then was successfully purged and has been declared an operational spare for the BZh-7 unit currently installed in the Elektron.  If used, BZh-5 will still use the standard internal micropumps.]

Before breakfast and physical exercise, CDR Padalka and FE/SO Fincke completed another session of the Russian crew health-monitoring program's medical assessment MO-9/Biochemical Urinalysis.  The FE stowed the Urolux hardware afterwards.   [MO-9 is conducted regularly every 30 days (and also before and after EVAs) and is one of five nominal Russian medical tests adopted by NASA for US crewmembers for IMG PHS (Integrated Medical Group/Periodic Health Status) evaluation as part of the "PHS/Without Blood Labs" exam.  The analysis uses the sophisticated in-vitro diagnostic apparatus Urolux developed originally for the Mir program.  Afterwards, the data were entered in the medical equipment computer (MEC)'s special IFEP software (In-Flight Examination Program).]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Subsequently, Padalka performed the third and last part of the current MBI-8 Profilaktika ("countermeasures") fitness assessment series, first with the usual blood tests (to determine lactate and creatine kinase levels in the blood with the AccuSport equipment), then by a physical exercise session on the TVIS treadmill, supported by tagup with a ground specialist.  Mike Fincke was available to assist as CMO (crew medical officer) as required.  (Last time done: 7/14)   [The TVIS test is identical to the MO-3 test performed on the treadmill in idling (non-motorized) mode with free choice of speeds within certain specified ranges (idle/walk/slow run/moderate run/fast run/walk/recovery).  In addition to the nominal test procedure, MBI-8/Part 3 calls for the use of the TEEM-100M gas analyzer during the test, the blood lactate measurements, and subjective evaluation of physical exertion levels (using the Borg Perceived Exertion Scale, viz., 10 steps from very light over hard and very hard to maximum) during the test.  At the end of the creatine kinase tests, the results were logged, copied from Cardiocassette-2000 recording to OCA for downlink, and reported to the ground via tagup.  The activity was also photo-documented with the Nikon D1 digital camera.]

The crew completed the periodic station air sampling.   [Mike Fincke first collected air samples with the GSC (grab sample container, #1004)) at the center of the SM, then used the new Dual Sorbent Tube (DST), instead of the old SSAS (Solid Sorbent Air Sampler), to gather samples in the center of the Lab and SM.  The CDR later took readings for air constituents in SM and FGB with the AK-1M sampler, then used the AK-1M-F sampler to test the SM and FGB air for Freon.  In addition, checking for CO (carbon monoxide) in the SM, Gennady used the IPD Draeger tubes gear.]

After the ground had temporarily inhibited the FDS (Fire Detection System), the crew, at ~9:15am EDT, began their second fire drill/OBT (on-board training), a mandatory periodic one-hour exercise specifically written for the current two-person crew.  Primary goal of this exercise is to provide the station residents with the most realistic emergency training possible.  The drill is always conducted with the support of both MCCs in close coordination.  (Last time done: 7/9).   [OBT objectives are to (a) practice fire response procedures (FRPs) and all incorporated actions for the case of a software-detected fire to locate, extinguish, and verify extinguishing attempts; (b) browse through RS laptop and the "Signal-VM" fire detection system displays as well as the automated software (algorithms) response to the fire event; (c) practice crew communication necessary to perform emergency FRPs;  (d) update the locations of support hardware (IPK-1M gas masks and OKR-1 fire extinguishers ) to be used for fire suppression in the RS.  The exercise does not actually use any fire equipment but simulates such actions to the maximum extent possible.  After the OBT, a post-training debrief was to be prepared.]

After the ground had activated the ER2 HRF (Human Research Facility) earlier in the morning, Mike powered up its laptop and set up the ADUM equipment, after which the crewmembers performed ultrasound bone scans (Scan Z) on each other, taking turns as operator and subject.  Afterwards the hardware was deactivated.  The scan heads were cleaned and stowed as part of closeout operations.   [The bone scans were taken of the subjects shoulder, elbow, knee and ankle, medically supported from the ground (POIC, Payload Operations & Integration Center) via privatized video and VOX S/G.  The data were also recorded, and the scanning and post-scan activities were videotaped and still-photographed for downlink.]  

Scheduled video inspection of IMV (intermodular ventilation) air ducts by Padalka, documenting with the DVCAM-150 camcorder, was deferred.

As a further step in the long-enduring struggle to recover the #2 air conditioner (SKV-2) of the Russian TCS (thermal control system) in the SM, Padalka installed components of a new replaceable condensate removal line (SMOK), essentially a flexhose with a T-adapter and an inlet fitting at the NOK condensate removal pump.   [The lingering problem on the SKV-2 had been identified on 6/22 as a leaky condensate removal line (air getting in through cracks and impeding condensate flow).  The new separate hose bypasses the plumbing between the SKV and the automatic NOK pump.]

The CDR terminated the bake-out cycle on the BMP micropurification system's channel #2, moding the channel back to Purify.  After yesterday's termination of regeneration on channel #1, this restored both filter beds to Purification/Absorption mode.  [The regeneration of the air purifier filter beds is repeated every 20 days.  Each bakeout to space vacuum takes about 24 hours.]  

The crew completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with bungee cord load trainer, and Mike performed the periodic (every other week) inspection of the RED.  

For the RED exercise, Mike Fincke set up the video equipment to record the session, providing documentation for ground specialists.  Afterwards, the video equipment was stowed again.   [The video is required for biomechanical evaluation of the exercising crewmember and assessment of the on-orbit setup of equipment during data collection.]  

Gennady conducted the periodic inspection of the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.  He also transferred the accumulated data files to storage for downlink to the ground via Regul-Paket (the Russian Email mode).

The FE completed the routine inspection of the SM's SOZh environment control & life support system and prepared the regular IMS (inventory management system) delta file for automated export to the ground for database updating and subsequent re-import via OCA.  

The OCA (orbital communications adapter) router laptop had problems booting up today and the screen was unreadable.  The shell of the laptop serving as the router was thought to be suspect.  The crew removed the router hard drive from the bad shell and placed it in the shell of SSC1 (space station computer 1).  The OCA is now operational.  

Gennady held his weekly IMS tagup with ground specialists, discussing open issues concerning identification of equipment and storage locations for updating the IMS database.   [Today's topics focused mainly on identifying and confirming stowage locations for EVA gear.]

Later tonight, FE/SO Fincke will unstow and install the equipment for the U.S. PHS/PCBA (periodic health status/portable clinical blood analyzer) with blood labs exam.  He also prepares the test equipment for the periodic Russian MO-10 "Hematokrit" testing.  Both sessions are scheduled for tomorrow.  [While PCBA analyzes total blood composition, MO-10 particularly measures the hematocrit (red blood cell mass) value of the blood (it is a well-known phenomenon of space flight that red blood cell mass {normal range: 30-45%} tends to go down over time).  

At 4:35pm, the crew is scheduled to have their regular (once every two weeks) teleconference with the JSC Astronaut Office (Kent Rominger), via S-band S/G.  

The crew was briefed on the current status and plans for the three on-board EMU (extravehicular mobility units) spacesuits.   [Of the three suits, two are functional, viz. EMU #3011 (although not tested for quite some time) and #3013 (with cooling restored by Mike recently, but not certified yet for use).  EMU #3005 has no cooling.  As discovered during the work on #3013, microbial growth in the cooling loop can contaminate the unit and jam the coolant pump.  Also, there were crystalline deposits (salts, etc.) in the heat exchanger, which can obstruct Airlock water lines.  On #3013, the pump was cleaned and a new rotor/impeller plus pump inlet filter installed.  For the rest of Increment 9, additional work on EMU #3005 will be prioritized to (a) perform a gas trap and pump inlet filter R&R, and (b) to determine the rate of microbial growth and other contaminants in the Airlock water lines by collecting appropriate water samples and flushing one of the cooling lines, taking a sample just before and immediately after the flushing.]  

Discussions were held at TsUP/Moscow and RSC-Energia on the upcoming ISS port relocation for the docked Soyuz spacecraft.  Of the two options, viz. letting relocation between DC1 "Pirs" and FGB nadir port be conducted by the Increment 9 crew (DC1-to-FGB, as proposed by Padalka) or the fresh Increment 10 crew (FGB-to-DC1, as in the original plan worked out jointly between Moscow and Houston) RSC-E head Yuri Semyonov decided in favor of the latter.  

Today's CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Beni River dynamics, Bolivia (this pass provided a good opportunity for oblique views of the river to the left and right of track.  Oblique views, used in conjunction with high-resolution nadir photography, are useful in mapping the past and present river channels), Patagonian Glaciers, South America (cloud cover was predicted to be relatively sparse at the time of ISS overpass.  Looking to the left of track; the sun position should have made for dramatic views of the mountains and glaciers), and Internal waves, Patagonian Shelf (a predicted clear weather window over Patagonia should have provided an opportunity to photograph internal waves off the coastline. Looking to the left of track for the sunglint point).

CEO images can be viewed at these websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 9 crew visit:

Major upcoming events:

  • Reboost -- 9/22 (phase angle correction for 9S)
  • Soyuz 9S launch -- 10/9, 1:04am EDT (w/Leroy Chiao, Salizhan Sharipov, Yuri Shargin);
  • Soyuz 9S dock -- 10/11, 2:15am EDT;
  • Soyuz 8S undock -- 10/18, 6:58pm EDT (w/Gennady Padalka, Michael Fincke, Yuri Shargin);
  • Soyuz 8S land -- 10/18, 10:22pm EDT;
  • Soyuz 9S port relocate -- 11/18;
  • Progress 15P undock – 11/23;
  • Progress 16P launch -- 11/24;
  • EVA-12 – 12/28;
  • Progress 16P undock – 1/29/05;
  • Progress 17P launch – 1/30/05;
  • EVA-13 – 2/21/05;
  • Shuttle/LF1 launch – NET 3/6/05;
  • Shuttle/LF1 undock – NET 3/16/05.

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

U.S. & Russian Segment Status  (as of today, 1:26pm EDT)   Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is Off.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is On.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is Off.  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On. SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) -- 739; temperature (deg C) -- 25.0; ppO2 (mmHg) -- 150.7; ppCO2 (mmHg) -- 3.2.
  • SM Transfer Compartment:  Pressure (mmHg) -- 741; temperature (deg C) -- 19.8.
  • FGB Cabin:  Pressure (mmHg) -- 744; temperature (deg C) -- 23.7.
  • Node:  Pressure (mmHg) -- 737.5; temperature (deg C) -- 24.6 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • U.S. Lab:  Pressure (mmHg) -- 739.4; temperature (deg C) -- 23.8; ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) -- 739.6; temperature (deg C) -- 28.4; shell heater temp (deg C) -- n/a, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
  • (n/a = data not available)

  Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, "sun slicer", i.e., drag reduction-biased by 47 deg angle (2B: +47, 4B: -47).
  • SM batteries:  Battery #1 is in "Cycle" mode; all other batteries (7) are on line in "Partial Charge" mode.  
  • FGB batteries:  Battery #3 is off line.  All other batteries (5) are on line in "Partial Charge" mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

  Command & Data Handling Systems (C&DH)

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

  Propulsion System (PS):

  • Total propellant load available: 4408 kg (9718 lb) as of 9/8/04;  [SM(552) + FGB(3388) + Progress M(468)].  (Capabilities: SM -- 860 kg; FGB -- 6120 kg).

  Attitude Control Systems (ACS):

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2's RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source -- U.S. SIGI-1 (GPS)
  • Attitude source -- U.S. SIGI-1 (GPS)
  • Angular rate source -- RGA-1

  Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = "sun-fixed" [yaw: 180.5 deg, pitch: -6.9 deg., roll: 0 deg]), with CMG TA (thruster assist) until 9/2 for EVA-11, then back to XPOP until next reboost (9/22).

  Communications & Tracking Systems (C&T):

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.


ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of this morning, 7:54am EDT [= epoch]):

  • Mean altitude -- 360.4 km
  • Apogee height -- 364.2 km
  • Perigee height -- 356.6 km
  • Period -- 91.75 min.
  • Inclination (to Equator) -- 51.63 deg
  • Eccentricity -- 0.0005614
  • Solar Beta Angle -- 56.7 deg (magnitude increasing)
  • Orbits per 24-hr. day -- 15.69
  • Mean altitude loss in last 24 hours -- 90 m
  • Revolutions since FGB/Zarya launch (Nov. 98)  -- 33156


ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA's Human Spaceflight website. The current location of the International Space Station can be found at at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.